a.) Find z0 such that P(−z0 < z < z0) = 0.6. (Round your answer to two decimal places.)

z0 =

b.) What percentile does −z0 represent? (Round your answer to the nearest whole number.)

c.) What percentile does z0 represent? (Round your answer to the nearest whole number.)
Answer #1

a) P (ozolz azo)= 0.6 P(zazo) – P (za-zo) = 6.6 P (zczo) - [1-P (ZLzo)] = 0.6 2P ( Zazo) -1 = ont 2 P (z czo) = 1.6 PLZCZo)=